OpenStudy (anonymous):

Solve the equation. 12r = 8

OpenStudy (anonymous):

r=2/3

OpenStudy (anonymous):

Simply bring the 12 over on the eight side. 12r = 8 r = 8/12 (simplfy by dividing by 4) r= 2/3